Exercise 15-9

Sheffield Corporation has 9,500 shares of $100 par value, 8%, preferred stock and 47,700 shares of $10 par value common stock outstanding at December 31, 2017.

Answer the questions in each of the following independent situations.

(a) If the preferred stock is cumulative and dividends were last paid on the preferred stock on December 31, 2014, what are the dividends in arrears at December 31, 2017?
Amount of dividends in arrears $


How should these dividends be reported?
The cumulative dividend is

reportednot reported

as a liability.

(b) If the preferred stock is convertible into 6 shares of $10 par value common stock and 4,400 shares are converted, what entry is required for the conversion assuming the preferred stock was issued at par value? (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit


(c) If the preferred stock was issued at $107 per share, how should the preferred stock be reported in the stockholders’ equity section? (Enter account name only and do not provide descriptive information.)

Sheffield Corporation
Balance Sheet (Partial)

Current AssetsCurrent LiabilitiesIntangible AssetsLong-term InvestmentsLong-term LiabilitiesPaid-in CapitalProperty, Plant and EquipmentStockholders' EquityTotal AssetsTotal Current AssetsTotal Current LiabilitiesTotal Intangible AssetsTotal LiabilitiesTotal Liabilities and Stockholders' EquityTotal Long-term InvestmentsTotal Long-term LiabilitiesTotal Paid-in CapitalTotal Property, Plant and EquipmentTotal Stockholders' Equity

$
Answer #1
Answer

(a) Annual Pref Share dividend = 8%*$100*9500 = $76000

So if Pref share dividend was last paid in De 2014, it is now due for 3 yrs till Dec2017.
So Total Cumulative dvidend due is 3* $76000 = $228000
(b) Each Pref Share will get converted to 6 Common Stock
So 4400 Pref Shares will result in 4400*6 = 26400 Common Stock
4400 Pref Shares of $100 each = $440,000
26400 common stock of $10 = $2,64,000
So Premium is $440000-$264,000 = $176,000 which is charged to Retained earnings
Journal entry will be
Convertible Preferred Stock Dr 440,000 $      4,40,000
Common Stock Cr 344,000 $       2,64,000
Retained Earnings Cr 86,000 $       1,76,000
(c)
Preferred stock, $100 par, 8%, 9500 shares authorized,
9500 shares issued and outstanding                                                  =$ 1,090,000
Contributed capital in excess of par value, Pref stock(7*9500)         = $ 66,500
Please Like ( positive Rating )